Week of 200511

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local:
  • remote: Andrew (TRIUMF), Miro (Chair, DB), Christoph (CMS), Maarten (ALICE), Vladimir (LHCb), Xin (BNL), Xavier (KIT), Jose (PIC), David (IN2P3), Dave (FNAL), Darren (RAL), Onno (NL-T1), Alberto (Compute), Vincent (Security), Olga (Batch), Ivan (ATLAS), Andrew (NL-T1), Julia (WLCG), Borja (Monitoring), Cristi (Storage)

Experiments round table:

  • CMS reports ( raw view) -
    • Production system loaded with many requests
      • Hitting some internal limits regarding request meta data handling
      • Less CPU usage overall

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Ongoing WG MC productions
    • Issues:
      • GRIDKA: Some ongoing file access issues. FZK are waiting on vendor. (GGUS:146379)

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • CNAF: NC
  • EGI: NC
  • FNAL: NTR
  • IN2P3: IN2P3-CC will be in maintenance on June 16th, a Tuesday. As usual details will be available one week before the event.
  • JINR: NTR
  • KISTI: NC
  • KIT: On yesterday evening, GPFS crashed for a yet unknown reason. Consequently, all SEs also stopped working, until our on-call engineer and storage experts revived GPFS.
  • NDGF: NTR
  • NL-T1: SARA-MATRIX dCache downtime on May 19th: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=28736. This minor dCache upgrade will bring us to the latest 6.1 release (which is ahead of the Golden Release schedule).
  • NRC-KI: NC
  • OSG: NC
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • Cooling issue at p8 ( OTG:0056418 ) - hoping cool weather continues until it is resolved, else potential capacity loss.
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Draft reports for the April 2020 availability sent around
  • MW Officer: NTR
  • Networks: NTR
  • Security: Recent security incident, see the details in the security broadcast sent on May 5th

AOB:

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2020-05-18 - MiroslavPotocky
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback